OpenStudy (anonymous):

How to simplify? (n+1)!/n!

OpenStudy (anonymous):

(n+1)! you can rewrite as (n+1)n!

OpenStudy (anonymous):

So the answer is (n+1)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So for (n+2)!/(n+3)!, the answer is just (n+2)/(n+3)?

OpenStudy (anonymous):

um no what is (n+3)! equal to ?

OpenStudy (anonymous):

(n+3)n! ??

OpenStudy (anonymous):

(n+3)!=(n+3)(n+2)(n+1)n.. and so on

OpenStudy (anonymous):

so you can rewrite that as (n+3)!=(n+3)(n+2)!

OpenStudy (anonymous):

Wait, so (n+2)! equals (n+2)(n+1) ?

OpenStudy (anonymous):

that is (n+2)(n+1)(n+0)(n-1)... and can go on , depends on what you have for n , lets say if we let n=1 we would have (1+2)(1+1)(1+0)(1-1) and that is because we have reached 0

OpenStudy (anonymous):

oh wait we dont include the 0 until we reach 1 my mistake sorry

OpenStudy (anonymous):

The problem says that it starts at n=1, so then that means that the answer for the second one is (n+3), right?

OpenStudy (anonymous):

ok so for the second one we have \[\frac{ (n+2)! }{ (n+3)! }=\frac{ (n+2)! }{ (n+3)(n+2)! }=\frac{ 1 }{ (n+3) }\]

OpenStudy (anonymous):

Ah, ok! I see!

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

no problem :)